OpenStudy (anonymous):

The cost of one pen is given by the expression. The cost of one pen is undefined when x=?

OpenStudy (anonymous):



OpenStudy (anonymous):

A. 3 and -2 B. -3 and 2 C. 3

OpenStudy (anonymous):

are you dividing

OpenStudy (anonymous):

\[2x^2-x-10 \ = \ 2x^2 +4x -5x -10 \ = \ 2x(x+2) -5(x+2) \ = \ (2x-5)(x+2)\] okay? I factorized the numerator polynomial.

OpenStudy (anonymous):

if so then here is the answer